I have three cousins on my dad's side. The product of their ages is 84. Next year the youngest will be 1/4 the age of the middle cousin. In 2 years from now, the oldest will be twice the current age of the middle. What is the age of my older cousin?
Answer by dfrazzetto(283)

cousins = x,y,z
xyz = 84
x+1 = 1/4(y+1)
z+2 = 2y

z = 2y - 2
x = y/4 - 3/4
plug into first equation;
(y/4 - 3/4)(y)(2y-2) = 84
y^2/4 - 3y/4 (2y - 2) = 84
((y^2 - 3y) / 4) (2y - 2) = 84
2y^3 - 2y^2 - 6y^2 + 6y = 84 * 4
2y^3 - 8y^2 + 6y = 336
y = 7
x = 1
z = 12